OpenStudy (holsteremission):

Consider the unit disk \(\mathcal{R}=\{(x,y)|x^2+y^2=1\}\). Let's say we partition it into \(p\) non-intersecting segments using \(p-1\) vertical lines such that the area of each segment is \(\dfrac{\pi}{p}\). Specifically, we're partitioning the interval \([-1,1]\) using a sequence \(x_{p,k}\) indexed by \(k\in\{0,1,2,\ldots,p\}\) so that \[-1=x_{p,0}<\cdots

OpenStudy (holsteremission):

In the case when \(p=3\), the partition is made up of the points \[\begin{cases}x_{3,0}=-1\\ x_{3,1}\approx -0.2649\\ x_{3,2}\approx 0.2649\\ x_{3,3}=1\end{cases}\] When \(p=4\), we have \[\begin{cases}x_{4,0}=-1\\ x_{4,1}\approx -0.404\\ x_{4,2}=0\\ x_{4,3}\approx 0.404\\ x_{4,4}=1\end{cases}\] And so on. The task is to find out whether or not the sum \(\displaystyle\sum_{n=1}^\infty x_{n,n-1}\) converges. If this diverges, what about \(\displaystyle\sum_{n=1}^\infty (-1)^{n-1}x_{n,n-1}\)? Bonus points for exact values if either case converges.

OpenStudy (holsteremission):

Traditional calc and numerical methods get me so far as determining each of the intermediate \(x_{p,k}\) values for a given. We only care about the rightmost one that's not \(1\), which is given by the solution to \[\arccos x_{p,k-1}-x_{p,k-1}\sqrt{1-{x_{p,k-1}}^2}=\frac{\pi}{p}\]

OpenStudy (holsteremission):

*for a given \(p\)*

OpenStudy (holsteremission):

I'm pretty sure the non-alternating series diverges, since it seems \(x_{n,n-1}\to1\), but haven't proved anything along these lines.

OpenStudy (kainui):

OK I think I finally understand what it is you're trying to do, now I'm actually gonna start thinking about how to start solving it; seems like what you're doing is reasonable. My knee jerk reaction is to like use a comparison test to sorta turn the square root into some other nice exponent or compare it to an integral which might be easier to evaluate. I really have no clue yet though what'll work but seems like a fun problem.

OpenStudy (holsteremission):

I'm glad it's understandable! I had a hell of a time formulating the problem with all these indices.

OpenStudy (kainui):

You could put an upper bound and lower bound on the circle by squeezing it between two trapezoids, it might somehow make it easier to prove that it converges/diverges that way too.

OpenStudy (kainui):

Err not areas, I guess you're doing like little arc lengths or something... Haha maybe I haven't quite understood it after all, it's somehow that second to last bit of all the circles with \(\pi/p\) area slices #_# wow this is confusing haha

OpenStudy (kainui):

I see now, it's the lengths of the projection of all the first cuts on the x-axis added up.

OpenStudy (holsteremission):

Right, the summand refers to the \(x\) coordinates of the dots in this quick sketch: